# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ V Towards a Quantum Optimization Utility
Figure 3: Optimal time $t^{*}_{\textrm{opt}}$ in seconds, normalized by the number of variables, to achieve a target performance $\alpha$ on a problem instance by problem instance basis. We consider simulated annealing and the Burer-Monteiro solvers on large problem sizes. The solid line $1-\alpha\sim(t^{*}_{\textrm{opt}}/N)^{-1/2}$ is a guide to the eye. The vertical dashed lines indicate the estimated runtime of the QRR* algorithm with $p=1$, $2$, and $3$ layers using $10^{3}$ qubits with $t_{\texttt{Rx}}=20$ ns, $t_{\texttt{ISWAP}}=50$ ns, $t_{\texttt{m+r}}=6~{}\mu$s, and a $p$-dependent optimal shot count [30]. Each data point is averaged over $1,000$ randomly generated problem instances for each number of variables. Error bars indicate one standard deviation.
@@
We now establish runtime and performance goals to assess the potential quantum utility of QRR* for the maximum cut problem on random $3$-regular graphs. Using the same methodology developed for Fig. 2, we estimate the optimal time $t^{*}_{\textrm{opt}}$ for the classical solvers SA and BM to achieve fictitious performance targets.
@@
With the duration of a step proportional to the number of edges in the input problem, the rescaled variable $k=t^{*}_{\textrm{opt}}/N$ relates directly to the number of steps for solving $3$-regular graphs. We empirically observe in Fig. 3 that in the regime considered, the average performance shows an asymptotic convergence that is consistent with $\sim k^{-1/2}$ for both SA and BM, independently of $N$. This power law in the context of SA is reminiscent of a Kibble-Zurek-type mechanism [39, 40], albeit different from the idealized statistical physics scenario [41, 42, 43, 44].
@@
In the regime $p\ll\ln N$, the runtime of QRR* is proportional to $N$ and moves towards $O(N^{2})$ for $p\gg\ln N$, as the QAOA starts to generate global quantum coherence across the whole graph [45, 30]. Hence, in the low-$p$ regime, QRR* has a linear algorithmic complexity with $N$ for a fixed performance, like its classical competitors. We display the runtime of QRR* rescaled by $N$ in Fig. 3 where an $M=10^{3}$ qubit quantum computer is assumed with $t_{\texttt{Rx}}=20$ ns and $t_{\texttt{ISWAP}}=50$ ns [30]. We observe that increasing $p$ substantially increases the runtime per variable because the light-cone technique yields circuits with sizes growing exponentially with $p$. Consequently, exponentially fewer such circuits can fit in parallel on $M$ qubits. The $p$-dependent runtime of QRR*, together with the scaling of classical solvers shown in Fig. 3, set performance targets for QRR* at fixed $p$ to achieve a quantum advantage for the maximum cut problem on $3$-regular graphs. For example, the QRR* algorithm with $p=1$ achieves $99\%$ approximation ratio (Fig. 1a). Under the assumptions of Fig. 3, QRR* should yield a fivefold average performance increase ($1-\alpha=1\%\to 0.2\%$) to be competitive in runtime with the SA and BM algorithms. Moving forward, while the exponential runtime increase with $p$ seems daunting, it might be counterbalanced by an exponential convergence of the quantum algorithm with $p$, as was observed for the QAOA [46] where $1-\alpha\sim e^{-p}$, albeit on problem sizes up to two orders of magnitude smaller than what we consider here, with the fate of larger problems remaining an open question.
VI Conclusion and Outlook
@@
Our results demonstrate the potential for quantum optimization to deliver high-quality solutions at speed on large-scale problems, comparable with state-of-the-art classical solvers. By leveraging the sparsity of 3-regular graphs, the light-cone technique made it possible to address problem sizes beyond what current quantum machines can naively handle. We developed QRR*, a high-performant hybrid quantum-classical algorithm inspired by classical semidefinite programming methods [27]. It finds optimal or near-optimal solutions with a single layer ($p=1$) of quantum logical operations and converges asymptotically to the optimal solution by increasing $p$ [27].
@@
The combined runtime of the quantum and classical components of QRR* is of the order of tenths of seconds on problems with thousands of variables [30]. This makes the approach competitive against Gurobi but short of others, such as simulated annealing and the Burer-Monteiro algorithms in the low-$p$ regime considered. As presented, it is unclear whether the QRR* framework could lead to a quantum advantage given the narrow path offered by increasing $p$ in a runtime-versus-performance tradeoff. To overcome these challenges, one direction is to develop better and more efficient hybrid quantum-classical solvers [5]. Another direction—not orthogonal to the other—is to investigate problems other than the maximum cut on $3$-regular graphs, which would potentially be more challenging for classical solvers and ideally easier for quantum ones. However, such problems are likely denser, rendering the light-cone technique inapplicable. It represents a challenge for superconducting platforms to encode a graph problem structure in a gate-based fashion with shallow depth to mitigate the noise limitations of contemporary hardware. It would be interesting to investigate the runtime of quantum error correction [47] overhead and what it means for delivering a practical quantum speedup against state-of-the-art classical optimization solvers.
@@ Acknowledgements.
We gratefully acknowledge L. T. Brady, N. Didier, S. Ejtemaee, S. Grabbe, S. Hadfield, P. A. Lott, F. B. Maciejewski, M. J. Reagor, E. G. Rieffel, D. Venturelli, Z. Wang, and T. Wilson for discussions, input, and collaborations on related works. This work is supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112090058 and IAA 8839, Annex 114. Authors from USRA also acknowledge support under NASA Academic Mission Services under contract No. NNA16BD14C. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards DDR-ERCAP0024427 and ASCR-ERCAP0028951. The experiments were performed through Rigetti Computing Inc.’s Quantum Cloud Services QCSTM on the AnkaaTM-2 superconducting quantum processor developed, fabricated, and operated by Rigetti Computing Inc.
@@
M. Dupont conceived and led the project with support from B. Sundar. D. E. Bernal Neira and Z. Peng performed the numerical simulations with the classical solver Gurobi. M. J. Hodson performed the runtime analysis of the quantum computer. B. Evert, M. J. Hodson, and S. Jeffrey developed software supporting experimental runs with randomized compilation and readout error mitigation. M. Dupont and B. Evert performed the experiments and collected the data. M. Dupont carried out the rest of the simulations and analyses. M. Dupont wrote the manuscript with contributions from all the co-authors. All co-authors contributed to the discussions leading to the completion of this project.
@@
M. Dupont, B. Evert, M. Hodson, S. Jeffrey, and B. Sundar are, have been, or may in the future be participants in incentive stock plans at Rigetti Computing Inc. M. Dupont and B. Sundar are inventors on two pending patent applications related to this work (No. 63/631,643 and No. 63/632,079). The other authors declare that they have no competing interests.
@@
Supplemental Materials for
Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer”
S1 Superconducting Quantum Processing Unit
S1.1 Hardware Characteristics
We performed experiments on the $84$-qubit superconducting quantum chip Rigetti AnkaaTM-2 based on a square lattice topology with tunable couplers. We only used a subset of the device consisting of a continuous linear chain of seven qubits with characteristics reported in Tab. S1. The device features parametric one-qubit gates, implementing rotations around the $x$ and $z$ axes, $\texttt{Rx}(\phi)=\exp(-i\hat{X}\phi/2)$ with $\phi=\pm\pi/2$, and $\texttt{Rz}(\phi\in\mathbb{R})=\exp(-i\hat{Z}\phi/2)$, respectively. The two-qubit gate between adjacent qubits is the ISWAP gate defined as $\exp[i\pi(\hat{X}\otimes\hat{X}+\hat{Y}\otimes\hat{Y})/4]$. $\hat{X}$, $\hat{Y}$, and $\hat{Z}$ are Pauli operators. We discuss the compilation of relevant quantum logical gates using natives ones in Sec. S4.1.2.
@@
Table S1: Characteristics of a linear chain of seven qubits used for running the experiments on the Rigetti AnkaaTM-2 superconducting chip. Reported fidelities were estimated using randomized benchmarking [50].
S1.2 Quantum Computer’s Output
Figure S1: Expectation value of the two-point correlation function $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ for two qubits $i$ and $j$ relevant for building the correlation matrices of all the graph problems based on the QAOA with $p=1$. Expectation values were estimated from $10^{4}$ bit strings (shots). There are $44$ entries corresponding to $44$ unique subgraphs according to Tab. S3. The data include noiseless state vector emulations and experimental data from the quantum computer using readout-error mitigation
We compare the emulated and device results for problems solved using QAOA with $p=1$. As shown in Sections S3.2 and S3.3, by leveraging the light cone induced by the QAOA with $p$ layers and graph isomorphism, we only need to evaluate a total of $44$ expectation values from $44$ different quantum circuits at $p=1$. The corresponding expectation values using readout error mitigation [34] (see also Sec. S4.1.3) are displayed in Fig. S1 with a clear correlation between emulated and simulated data. The overall magnitude reduction of the experimental data is expected and understood under a simple depolarizing noise model [27], as developed in Sec. S4.1.3.
We note that graph isomorphism is just a practical convenience that enables us to execute a fixed number of circuits at fixed $p$ to get statistics over any number of problem instances. Its use is not assumed when estimating the runtime of the quantum algorithm.
S2 Combinatorial Optimization Solvers
@@
$\tilde{C}(\boldsymbol{z})=\frac{1}{2}\sum\nolimits_{i,j=1}^{N}\mathsf{W}_{ij}z_{i}z_{j},$ (S1)
where $\mathsf{W}\in\mathbb{R}^{N\times N}$ is a symmetric matrix, i.e., $\mathsf{W}_{ij}=\mathsf{W}_{ji}$, encoding the problem of interest. The variables $z_{i}$ are subject to a binary constraint such that $\boldsymbol{z}\in\{\pm 1\}^{N}$. Specifically, we focus on minimizing the objective function of Eq. (S1) where $\mathsf{W}$ is drawn uniformly from the ensemble of adjacency matrices of unit-weight $3$-regular graphs. As such, $\mathsf{W}$ is sparse with three nonzero entries per row and column, each equal to unity.
Up to an additional factor and overall multiplicative constant, minimizing Eq. (S1) is similar to maximizing the objective function of the main text for the maximum cut problem. When needed, we characterize the quality of a solution $\boldsymbol{z}$ by the approximation ratio
$\alpha=\frac{C(\boldsymbol{z})}{C(\boldsymbol{z}_{\textrm{opt}})}=\frac{3N-2\tilde{C}(\boldsymbol{z})}{3N-2\tilde{C}(\boldsymbol{z}_{\textrm{opt}})},$ (S2)
where $C$ is the maximum cut objective function defined in the main text, $\tilde{C}$ the related objective function of Eq. (S1), and $\boldsymbol{z_{\textrm{opt}}}$ is the optimal solution.
S2.1 Classical
S2.1.1 Markov Chain Monte-Carlo
The objective function $\tilde{C}(\boldsymbol{z})$ of Eq. (S1) describes an antiferromagnetic Ising model. The optimal solution $\boldsymbol{z}_{\textrm{opt}}$, which minimizes the objective function, is the ground state of the Ising model. Thus, in a canonical ensemble, the optimal solution $\boldsymbol{z}_{\textrm{opt}}$ would perfectly describe the state of the model at zero temperature, i.e., $T=0$. More generally, at temperature $T$, the probability for the model to be in the state $\boldsymbol{z}$ with energy $\tilde{C}(\boldsymbol{z})$ is given by a Boltzmann distribution
$p(T,\boldsymbol{z})\propto e^{-\tilde{C}(\boldsymbol{z})/T}.$ (S3)
The Metropolis-Hastings Markov chain Monte-Carlo algorithm [51, 52] is a celebrated approach for numerically evaluating the thermodynamic properties of an Ising model at temperature $T$. It is based on the detailed balance principle, which samples microstates $\boldsymbol{z}$ according to Eq. (S3), after a thermalization process. From a configuration $\boldsymbol{z}$, a new configuration $\boldsymbol{z^{\prime}}$ is suggested. It is accepted with a Metropolis-Hastings probability
$p_{\textrm{accept}}\bigl{(}\boldsymbol{z}\leftarrow\boldsymbol{z^{\prime}}\bigr{)}=\textrm{min}\Biggl{\{}1,e^{\bigl{[}\tilde{C}(\boldsymbol{z})-\tilde{C}(\boldsymbol{z^{\prime}})\bigr{]}/T}\Biggr{\}}.$ (S4)

### v2

@@ V Towards a Quantum Optimization Utility
Figure 3: The optimal time $t^{*}_{\textrm{opt}}$, in seconds, normalized by the number of variables, to achieve a target performance $\alpha$ on a problem instance by problem-instance basis. We consider simulated annealing and the Burer-Monteiro solvers on large problem sizes. The solid line $1-\alpha\sim(t^{*}_{\textrm{opt}}/N)^{-1/2}$ is a guide to the eye. The vertical dashed lines indicate the estimated run-time of the QRR* algorithm with $p=1$, $2$, and $3$ layers, using $10^{3}$ qubits with $t_{\texttt{Rx}}=20$ ns, $t_{\texttt{ISWAP}}=50$ ns, $t_{\texttt{m+r}}=6~{}\mu$s, and a $p$-dependent optimal shot count (Appendix E.3). Each data point has been averaged over $1,000$ randomly generated problem instances for each number of variables. The error bars indicate one standard deviation.
@@
We now establish run-time and performance goals to assess the potential quantum utility of the QRR* for the maximum-cut problem on random $3$-regular graphs. Using the same methodology as developed for Fig. 2, we estimate the optimal time $t^{*}_{\textrm{opt}}$ for the classical SA and BM solvers to achieve fictitious performance targets.
@@
With the duration of a step proportional to the number of edges in the input problem, the rescaled variable $k=t^{*}_{\textrm{opt}}/N$ relates directly to the number of steps for solving $3$-regular graphs. We empirically observe in Fig. 3 that in the regime considered, the distance to the optimal solution is reduced asymptotically as approximately $k^{-1/2}$ for both SA and the BM, independently of $N$. This power law in the context of SA is reminiscent of a Kibble-Zurek-type mechanism [39, 40], albeit different from the idealized statistical-physics scenario [41, 42, 43, 44].
@@
In the regime $p\ll\ln N$, the run-time of the QRR* is proportional to $N$ and moves toward $O(N^{2})$ for $p\gg\ln N$, as the QAOA starts to generate global quantum coherence across the whole graph [45]. Hence, in the low-$p$ regime, the QRR* has a linear algorithmic complexity with $N$ for a fixed performance, like its classical competitors. We display the run-time of the QRR* rescaled by $N$ in Fig. 3, where an $M=10^{3}$ qubit quantum computer is assumed with $t_{\texttt{Rx}}=20$ ns and $t_{\texttt{ISWAP}}=50$ ns. We observe that increasing $p$ substantially increases the run-time per variable because the light-cone technique yields circuits with sizes growing exponentially with $p$. Consequently, exponentially fewer such circuits can fit in parallel on $M$ qubits. The $p$-dependent run-time of the QRR*, together with the scaling of classical solvers shown in Fig. 3, set performance targets for the QRR* at fixed $p$ to achieve a quantum advantage for the maximum-cut problem on $3$-regular graphs. For example, the QRR* algorithm with $p=1$ achieves $99\%$ approximation ratio (Fig. 1a). Under the assumptions of Fig. 3, QRR* should yield a fivefold average performance increase ($1-\alpha=1\%\to 0.2\%$) to be competitive in run-time with the SA and BM algorithms. Moving forward, while the exponential run-time increase with $p$ seems daunting, it might be counterbalanced by an exponential convergence of the quantum algorithm with $p$, as has been observed for the QAOA [46], where $1-\alpha\sim e^{-p}$, albeit on problem sizes up to $2$ orders of magnitude smaller than what we consider here, with the fate of larger problems remaining an open question.
VI Conclusions and Outlook
@@
Our results demonstrate the potential for quantum optimization to deliver high-quality solutions at speed on large-scale problems, comparable with state-of-the-art classical solvers. By leveraging the sparsity of 3-regular graphs, the light-cone technique has made it possible to address problem sizes beyond what current quantum machines can naively handle. We have developed the QRR*, a high-performant hybrid quantum-classical algorithm inspired by classical semidefinite programming methods [29]. It finds optimal or near-optimal solutions with a single layer ($p=1$) of quantum logical operations and converges asymptotically to the optimal solution by increasing $p$ [29].
@@
The combined run-time of the quantum and classical components of the QRR* is of the order of tenths of seconds on problems with thousands of variables (Appendix E.3). This makes the approach competitive against Gurobi but short of others, such as the SA and BM algorithms in the low-$p$ regime considered. As presented, it is unclear whether the QRR* framework could lead to a quantum advantage given the narrow path offered by increasing $p$ in a run-time-versus-performance tradeoff. To overcome these challenges, one direction is to develop better and more efficient hybrid quantum-classical solvers [5]. Another direction—not orthogonal to the other—is to investigate problems other than the maximum cut on $3$-regular graphs, which would potentially be more challenging for classical solvers and ideally easier for quantum ones. However, such problems are likely denser, rendering the light-cone technique inapplicable. It represents a challenge for superconducting platforms to encode a graph-problem structure in a gate-based fashion with shallow depth to mitigate the noise limitations of contemporary hardware. It would be interesting to investigate the run-time of quantum error correction [47] overhead and what it means for delivering a practical quantum speedup against state-of-the-art classical optimization solvers.
@@ Acknowledgements.
We gratefully acknowledge L. T. Brady, N. Didier, S. Ejtemaee, S. Grabbe, S. Hadfield, P. A. Lott, F. B. Maciejewski, M. J. Reagor, E. G. Rieffel, D. Venturelli, Z. Wang, and T. Wilson for discussions, input, and collaborations on related works. This work is supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112090058 and IAA 8839, Annex 114. The authors from USRA also acknowledge support by NASA Academic Mission Services under Contract No. NNA16BD14C. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy (DOE) Office of Science User Facility supported by the Office of Science of the DOE under Contract No. DE-AC02-05CH11231 using NERSC Awards DDR-ERCAP0024427 and No. ASCR-ERCAP0028951. The experiments were performed through Rigetti Computing Inc.’s Quantum Cloud Services QCSTM on the AnkaaTM-2 superconducting quantum processor developed, fabricated, and operated by Rigetti Computing Inc.
@@
M.D. conceived and led the project with support from B.S. D.E.B.N. and Z.P. performed the numerical simulations with the classical solver Gurobi. M.J.H. performed the run-time analysis of the quantum computer. B.E., M.J.H., and S.J. developed software supporting the experimental runs with randomized compilation and readout-error mitigation. M.D. and B.E. performed the experiments and collected the data. M.D. carried out the rest of the simulations and analyses. M.D. wrote the manuscript, with contributions from all the coauthors. All coauthors contributed to the discussions leading to the completion of this project.
@@
M.D., B.E., M.H., S.J., and B.S. are, have been, or may in the future be participants in incentive stock plans at Rigetti Computing Inc. M.D. and B.S. are inventors on two pending patent applications related to this work (No. 63/631,643 and No. 63/632,079). The other authors declare that they have no competing interests.
@@
Appendix A Superconducting Quantum Processing Unit
A.1 Hardware Characteristics
We performed experiments on the $84$-qubit superconducting quantum chip Rigetti AnkaaTM-2 based on a square lattice topology with tunable couplers. We only used a subset of the device consisting of a continuous linear chain of seven qubits with characteristics reported in Tab. 1. The device features parametric one-qubit gates, implementing rotations around the $x$ and $z$ axes, $\texttt{Rx}(\phi)=\exp(-i\hat{X}\phi/2)$ with $\phi=\pm\pi/2$, and $\texttt{Rz}(\phi\in\mathbb{R})=\exp(-i\hat{Z}\phi/2)$, respectively. The two-qubit gate between adjacent qubits is the ISWAP gate defined as $\exp[i\pi(\hat{X}\otimes\hat{X}+\hat{Y}\otimes\hat{Y})/4]$. $\hat{X}$, $\hat{Y}$, and $\hat{Z}$ are Pauli operators. We discuss the compilation of relevant quantum logical gates using natives ones in App. D.1.2.
@@ A.1 Hardware Characteristics
Table 1: Characteristics of a linear chain of seven qubits used for running the experiments on the Rigetti AnkaaTM-2 superconducting chip. Reported fidelities were estimated using randomized benchmarking [48].
A.2 Quantum Computer’s Output
Figure 4: Expectation value of the two-point correlation function $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ for two qubits $i$ and $j$ relevant for building the correlation matrices of all the graph problems based on the QAOA with $p=1$. Expectation values were estimated from $10^{4}$ bit strings (shots). There are $44$ entries corresponding to $44$ unique subgraphs according to Tab. 3. The data include noiseless state vector emulations and experimental data from the quantum computer using readout-error mitigation
We compare the emulated and device results for problems solved using QAOA with $p=1$. As shown in Sections C.2 and C.3, by leveraging the light cone induced by the QAOA with $p$ layers and graph isomorphism, we only need to evaluate a total of $44$ expectation values from $44$ different quantum circuits at $p=1$. The corresponding expectation values using readout error mitigation [36] (see also App. D.1.3) are displayed in Fig. 4 with a clear correlation between emulated and simulated data. The overall magnitude reduction of the experimental data is expected and understood under a simple depolarizing noise model [29], as developed in App. D.1.3.
We note that graph isomorphism is just a practical convenience that enables us to execute a fixed number of circuits at fixed $p$ to get statistics over any number of problem instances. Its use is not assumed when estimating the run-time of the quantum algorithm.
Appendix B Combinatorial Optimization Solvers
@@
$\tilde{C}(\boldsymbol{z})=\frac{1}{2}\sum\nolimits_{i,j=1}^{N}\mathsf{W}_{ij}z_{i}z_{j},$ (4)
where $\mathsf{W}\in\mathbb{R}^{N\times N}$ is a symmetric matrix, i.e., $\mathsf{W}_{ij}=\mathsf{W}_{ji}$, encoding the problem of interest. The variables $z_{i}$ are subject to a binary constraint such that $\boldsymbol{z}\in\{\pm 1\}^{N}$. Specifically, we focus on minimizing the objective function of Eq. (4) where $\mathsf{W}$ is drawn uniformly from the ensemble of adjacency matrices of unit-weight $3$-regular graphs. As such, $\mathsf{W}$ is sparse with three nonzero entries per row and column, each equal to unity.
Up to an additional factor and overall multiplicative constant, minimizing Eq. (4) is similar to maximizing the objective function of the main text for the maximum-cut problem. When needed, we characterize the quality of a solution $\boldsymbol{z}$ by the approximation ratio
$\alpha=\frac{C(\boldsymbol{z})}{C(\boldsymbol{z}_{\textrm{opt}})}=\frac{3N-2\tilde{C}(\boldsymbol{z})}{3N-2\tilde{C}(\boldsymbol{z}_{\textrm{opt}})},$ (5)
where $C$ is the maximum cut objective function defined in the main text, $\tilde{C}$ the related objective function of Eq. (4), and $\boldsymbol{z_{\textrm{opt}}}$ is the optimal solution.
B.1 Classical
B.1.1 Markov Chain Monte-Carlo
The objective function $\tilde{C}(\boldsymbol{z})$ of Eq. (4) describes an antiferromagnetic Ising model. The optimal solution $\boldsymbol{z}_{\textrm{opt}}$, which minimizes the objective function, is the ground state of the Ising model. Thus, in a canonical ensemble, the optimal solution $\boldsymbol{z}_{\textrm{opt}}$ would perfectly describe the state of the model at zero temperature, i.e., $T=0$. More generally, at temperature $T$, the probability for the model to be in the state $\boldsymbol{z}$ with energy $\tilde{C}(\boldsymbol{z})$ is given by a Boltzmann distribution
$p(T,\boldsymbol{z})\propto e^{-\tilde{C}(\boldsymbol{z})/T}.$ (6)
The Metropolis-Hastings Markov chain Monte-Carlo algorithm [49, 50] is a celebrated approach for numerically evaluating the thermodynamic properties of an Ising model at temperature $T$. It is based on the detailed balance principle, which samples microstates $\boldsymbol{z}$ according to Eq. (6), after a thermalization process. From a configuration $\boldsymbol{z}$, a new configuration $\boldsymbol{z^{\prime}}$ is suggested. It is accepted with a Metropolis-Hastings probability
$p_{\textrm{accept}}\bigl{(}\boldsymbol{z}\leftarrow\boldsymbol{z^{\prime}}\bigr{)}=\textrm{min}\Biggl{\{}1,e^{\bigl{[}\tilde{C}(\boldsymbol{z})-\tilde{C}(\boldsymbol{z^{\prime}})\bigr{]}/T}\Biggr{\}}.$ (7)